# Adaptive Partition Factor Analysis

### v1

@@ 1 Introduction
In numerous scientific and socio-economic domains, the collection of high-dimensional data has become ubiquitous. Instances of such data collection include customer preferences in recommender system applications, single-cell experiments in genomics, high-frequency trading in finance, and extensive sampling campaigns in ecology. In these contexts, the joint analysis of data originating from diverse sources, studies, or technologies is increasingly critical for enhancing the accuracy of conclusions and ensuring the reproducibility of results [26].
However, it is common that different studies mix the common latent signal with study-specific ones. This issue is particularly pronounced in our motivating contexts of animal co-occurrence studies and genomics. In the former, zoothologists record the presence of various animal species across different spatial locations, conducting sampling campaigns over multiple time periods. While most sites with similar habitats exhibit consistent patterns of animal occurrences, some sites may display unique characteristics due to possible unobserved variables [29]. Similarly, in gene expression studies, high-throughput experiments are known to present variations of both biological and technical nature. Consequently, effects driven by specific experimental conditions or particular technology can overshadow the biological signals of interest [27, 28]. Although preprocessing procedures aimed at correcting for batch effects help in removing some specificities, these methods are often not enough to obtain homogeneous covariance structures across groups [21].
Motivated by these problems, [1] introduced a novel methodological tool, namely multi-study factor analysis (MSFA), which extends the classical factor analysis to jointly analyze data from multiple studies. MSFA aims to separate the signal shared across multiple studies from the study-specific components arising from artifactual or population-specific sources of variation. To fix the notation, let $y_{is}$ denote the $p$-dimensional observation vector corresponding to subject $i\in\{1,\dots,n\}$ belonging to study $s\in\{1,\dots,S\}$, where each study comprises $n_{s}$ subjects and $n=\sum_{s=1}^{S}n_{s}$. In the MSFA framework, one assumes the following representation:
@@
where $\epsilon_{is}$ represents a zero-mean idiosyncratic error term and $\eta_{is}$ is a $d$ dimensional set of shared latent factors, with $d\ll p$ and with $\Lambda$ the corresponding factor loading matrix of dimension $p\times d$. Similarly, $\Gamma_{s}$ is a study-specific factor loading matrix of dimension $p\times k_{s}$, with $k_{s}\ll p$ possibly different in each study, and ${\varphi}_{is}$ its corresponding latent factors.
@@
To address the high-$p$ small-$n$ problem typical of high-throughput biological data, [8] propose a Bayesian generalization of [1] that naturally provides the necessary regularization through the multiplicative gamma prior of [23]. Posterior sampling is performed via Markov Chain Monte Carlo (MCMC) methods, enabling flexible estimation without constraints on loading matrices. Extensions handling covariates and alternative estimation procedures has been discussed by [20] and [21]. Similarly, [32] proposed a perturbed factor analysis that focuses on inferring the shared structure while making use of subject-specific perturbations. Motivated by identification issues arising from these approaches, [3] recently proposed a class of subspace factor models which characterize variation across groups at the level of a lower-dimensional subspace where the study-specific factor loadings $\Gamma_{s}$ are obtained as $\Gamma_{s}=\Lambda A_{s}$ where $A_{s}$ is a $d\times d_{s}$ matrix, with $d_{s}<d$, i.e. modeling the study-specific contributions as lower dimensional perturbations of the shared factor loadings.
@@
In all of these approaches there is however a modeling limitation, i.e. they allow precisely $S$ study-specific loading matrices $\Gamma_{s}$. In practice, this assumption is often violated. We provide some examples, but a continuous spectrum of possibilities may be found: (i) two or more studies exhibit high homogeneity and may share identical or nearly identical latent representations, and (ii) one study involves highly heterogeneous groups of subjects, potentially resulting in two or more sub-populations with distinct latent structures as described in Equation (1). The latter situation might be practically due to slightly different experimental setups or unmeasured confunders. Additional possibilities include (iii) the existence of a subset of units from different groups that share common characteristics, or (iv) instances where, for various reasons, some units may be associated with characteristics of multiple groups simultaneously.
@@
To address case (i), [19] proposed a model allowing for partially-shared latent factors using a study-specific diagonal selection matrix. The pattern of zeros and ones in this matrix determines which latent factors pertain to each specific study. The diagonal elements are then incorporated into a selection matrix governed by an Indian Buffet Process (IBP) prior [30]. The same work discusses a variant, where study-specific labels are unknown and inferred via mixture models. While this approach moves towards addressing case (ii), it relies on the a priori knowledge of the total number of groups into which the units are divided and to our knowledge it has not been applied yet. Case (iii) corresponds to a situation where instead a single partition is insufficient to describe specific factors patterns. Hence, despite these recent advancements, a unifying framework that seamlessly spans from classical factor models to MSFA, encompassing all the above mentioned cases or nuanced variants, remains absent.
@@
In this paper, we propose an alternative formulation for Bayesian MSFA leveraging the concept of informed or structured sparsity [14, 15, 31]. This approach allows for flexible modeling of varying degrees of group heterogeneity in terms of latent factors contribution. Consistently with this, we name the proposed approach Adaptive Partition Factor Analysis (APAFA). APAFA employs shrinkage priors for latent factors, using the information contained in study labels, but without enforcing a deterministic structure. This structured shrinkage is combined with increasing shrinkage priors on the number of latent factors, following a cumulative shrinkage process approach [16, 22]. This construction ensures identifiable separation of the shared and specific contributions while circumventing, under suitable regularity conditions, the well-known issue of rotational ambiguity in factor loadings for the specific factors, thereby allowing straightforward posterior interpretation of these quantities. Notably, we show that the proposed method can be seen as a particular neural network model allowing us to exploit further generalizations, including the integration of subject- or study-specific continuous covariates in a flexible manner.
@@
The next section details APAFA in terms of model specification, prior distributions, and provides practical guidelines for prior elicitation. We also explore the connections to neural networks and present generalizations. In addition, we address model identifiability and describe the procedures for posterior computation. Section 3 reports a comprehensive empirical assessment through simulation studies covering various data-generating scenarios. Section 4 reports two illustrative data analyses further highlighting the advantages of our approach in the motivating examples of bird species occurrence and genomics. The final discussion section explores generalizations and potential extensions. The code to reproduce our analyses available at https://github.com/elenabortolato/APAFA.
@@ 2.1 Model and prior specification
where $\Gamma=(\Gamma_{1},\dots,\Gamma_{S})$ stacks all the study-specific factor loading matrices into a $p\times k$ matrix with $k=\sum_{s=1}^{S}k_{s}$ and ${\varphi}_{i}$ is a $k$-dimensional augmented vector containing the ${\varphi}_{is}$ in Equation (1) framed with suitable pattern of zeros. Specifically, all the latent factors pertaining to group $s$ will have $\sum_{l=1}^{s-1}k_{l}$ zeros, followed by subject-specific ${\varphi}_{is}$, followed by $\sum_{l=s+1}^{S}k_{l}$ zeros. An illustration is depicted in Figure 1.
@@
Figure 1: Multi study Factor model: The $n\times p$ data matrix $Y$ is written as the product of the latent factor matrix $H$ of dimension $n\times d$ by the factor loading matrix $\Lambda$ (the shared parts) plus the product of the latent factor matrix $\Phi$ of dimension $n\times k$ by the factor loading matrix $\Gamma$ (the study-specific parts) and a random noise $\epsilon$. Different shades of purple, red, and blue identify the $S=3$ studies in $Y$. The matrix $\Phi$ presents a pronounced presence of structural zeros and study-specific factors.
As customary in factor analysis, one can marginalize out the shared factors, obtaining
@@
Under this characterization, one can think to have, for each $i$, a vector of dummy variables $x_{i}$ with $S$ entries characterizing the study to which unit $i$ belongs to and then assume that the $h$-th element of the vector $\varphi_{i}$ is $\varphi_{ih}=\tilde{\varphi}_{ih}\psi_{ih}$ where $\tilde{\varphi}_{ih}$ is a continuous random variable and $\psi_{ih}=f_{h}(x_{i})$ with $f_{h}$ a deterministic activation function. In [1, 8], for example, $f_{h}(x_{i})=x_{i}^{\top}1_{S}$ where $1_{S}$ is a $S$ dimensional vector of ones. In this paper, we propose to incorporate the information contained in the $x_{i}$’s in a more flexible manner letting
@@
It is now evident that equations (3) and (4) transform model (2) into a specific single layer neural network where the dummy variables in the $x_{i}$s are the input variables, $y_{i}$ are the $p-$dimensional output variables, $\varphi_{i}$s are the nodes of the hidden layer, $f_{h}$ are the activation functions, $\beta_{h}$ are the weights between the input and the hidden layer, and the elements in $\Gamma$ are the weights between the hidden layer and the continuous outcome $y_{i}$. The neural network representation of model (2) is reported in Figure 2. In recent years, there has been a growing interest in Bayesian analysis of neural networks. Typically, the weight parameters are assigned a prior distribution and are learned updating the prior via Bayes rule. For comprehensive reviews, see [38], [11], and [39]. Bayesian approaches to neural network models, however, not only apply the Bayesian inferential paradigm to learn the network parameters but also allow the neurons themselves to be stochastic [34, 35].
@@
We let the global scale parameter $\tau^{\varphi}_{h}\sim\text{Ber}(1-\rho_{h})$, where $\{\rho_{h}\}\sim\mbox{CUSP}(\alpha^{\varphi})$ follows a cumulative shrinkage process [16], i.e. for $h=1,2,\dots,$ $\rho_{h}=\sum_{l\leq h}w^{\varphi}_{l}$,   $w^{\varphi}_{l}=v^{\varphi}_{l}\prod_{m<l}(1-v^{\varphi}_{l})$, and $v^{\varphi}_{l}\sim\text{Beta}(1,\alpha^{\varphi})$. This approach facilitates learning the number of study-specific factors or, following the neural network interpretation of our model, the number of neurons in the hidden layer. Notably, this construction, inspired by successful applications in factor models [16, 14], offers valuable insights into performing Bayesian inference for the number of nodes in the hidden layer of a Bayesian neural network, potentially connecting with existing literature on this topic [e.g 9, 41, 11, 10, 13, 12].
@@
$\tilde{\varphi}_{ih}(x_{i})\sim N(0,\tau_{h}^{\varphi}).$ (7)
Marginalizing out $\psi_{ih}(x_{i})$ with $\tau_{h}^{\varphi}$ and $\beta_{h}$ fixed, we have
@@
where $\delta_{a}$ is a Dirac mass at value $a$. Marginalizing out $\tilde{\varphi}_{ih}$, having unitary variance leads to
$y_{i}\sim N(0,\Lambda\Lambda^{\top}+\Gamma\text{diag}{\{\psi_{i}\}}\Gamma^{\top}+\Sigma).$ (8)
Notably (6) corresponds to a simple form of conditional variational autoencoder that, differently from the standard, allows for the latent covariance structure to be informed by the independent variables. Variational autoencoders are specific extension of neural networks, widely recognized as a descendant of classical factor models [36, 37].
The benefits of the proposed solution should now be evident. The shrinkage prior on the elements $\varphi_{ih}(x_{i})$ enables and promotes, yet does not mandate, the sparse representation in (2). Furthermore, it accommodates a wide range of scenarios discussed in the Introduction, including: (i) two or more studies exhibit high homogeneity and share nearly identical latent representations, (ii) some studies involve highly heterogeneous groups of subjects, potentially resulting in two or more sub-populations with distinct latent structures, or any combinations of the above. The random partition of the observation induced by the sparse structure of the $\psi_{i}$ justify the Adaptive Partition Factor Analysis (APAFA) name adopted for the proposed solution.
@@
One remark on the activation function pertains the potential lack of properness of the conditional posterior distribution for the parameters $\beta_{h}$, $h=1,\dots,k$. This concern is tied to the issue of perfect separation in logistic regression. Assuming the group structure encoded by the $x_{i}$ dummy variables is both accurate and relevant to the problem at hand—in other words, that the original MSFA model by [8] is correctly specified– a perfect separation in the $\psi_{ih}$ occurs with respect to the $x_{i}$ within the logistic regression. For this reason, the prior distribution on each parameter $\beta_{h}$ should be sufficiently concentrated around zero to prevent the associated conditional posterior from exhibiting a tendency toward monotonicity. To this end, we specify $\beta_{h}\sim N(0,B)$, where $B$ is a $S\times S$ diagonal matrix with elements proportional to $n^{-1}$. Thus, when perfect separation in the $\psi_{ih}$’s with respect to the $x_{is}$’s occurs, the full conditional posterior is

### v2

@@ 1 Introduction
In numerous scientific and socio-economic domains, the collection of high-dimensional data has become ubiquitous. Instances of such data collection include customer preferences in recommender system applications, single-cell experiments in genomics, high-frequency trading in finance, and extensive sampling campaigns in ecology. In these contexts, joint analysis of data originating from diverse sources, studies, or technologies is increasingly critical for enhancing the accuracy of conclusions and ensuring the reproducibility of results (National Academies of Sciences et al., 2019).
In these contexts, factor analysis is a standard technique for identifying common patterns of variation in multivariate data and relating it to hidden causes, the so-called latent factors. However, a key challenge arises when the latent factors identified across different studies combine both shared factors (common across studies) and factors unique to individual studies. This issue is particularly pronounced in our motivating contexts of animal co-occurrence studies and genomics. For instance, species occurrence data collected across spatial-temporal surveys often exhibit site-specific deviations despite shared environmental conditions (Ovaskainen et al., 2017). Similarly, high-throughput genomic experiments are prone to technical variability, where artifacts from experimental conditions may dominate biological signals (Irizarry et al., 2003; Shi et al., 2006; Vallejos et al., 2017; Hicks et al., 2018). While preprocessing adjusts for some biases, residual heterogeneity in covariance structures often persists across groups, challenging assumptions of homogeneity (Avalos-Pacheco et al., 2022).
Motivated by these problems, De Vito et al. (2019) introduced a novel methodological tool, namely multi-study factor analysis (MSFA), which extends the classical factor analysis to jointly analyze data from multiple studies. MSFA aims to separate the signal shared across multiple studies from the study-specific components arising from artifactual or population-specific sources of variation. Let ${y}_{is}\in\mathbb{R}^{p}$ denote the $p$-dimensional observation vector for the $i$-th subject ($i=1,\dots,n_{s}$) in the $s$-th study ($s=1,\dots,S$), where each study $s$ has $n_{s}$ subjects and the total sample size is $n=\sum_{s=1}^{S}n_{s}$. In the MSFA framework, one assumes the following representation:
@@
where $\epsilon_{is}$ represents a zero-mean idiosyncratic error term and $\eta_{is}$ is a $d$-dimensional set of shared latent factors, with $d\ll p$ and with $\Lambda$ the corresponding factor loading matrix of dimension $p\times d$. Similarly, $\Gamma_{s}$ is a study-specific factor loading matrix of dimension $p\times k_{s}$, with $k_{s}\ll p$, and ${\varphi}_{is}$ its corresponding latent factors. Notably, each subject belongs to a single study and, in fact, the index $s$ in the notation $y_{is}$ may be redundant as also discussed later in Section 2. Figure 1 reports a pictorial representation of the data matrix $Y$ with $n$ rows defined as $Y=(y_{i1}^{\top},\dots,y_{nS}^{\top})^{\top}$.
@@
To address the high-$p$ small-$n$ problem typical of high-throughput biological data, De Vito et al. (2021) propose a Bayesian generalization of the MSFA model that naturally provides regularization through the multiplicative gamma prior of Bhattacharya & Dunson (2011). Posterior sampling is performed via Markov Chain Monte Carlo (MCMC) methods, enabling flexible estimation without constraints on loading matrices. Extensions handling covariates and alternative estimation procedures has been later discussed by De Vito & Avalos-Pacheco (2023) and Avalos-Pacheco et al. (2022). Similarly, Roy et al. (2021) proposed a perturbed factor analysis that focuses on inferring the shared structure while making use of subject-specific perturbations. Motivated by identification issues arising from these approaches, Chandra et al. (2024) recently proposed a class of subspace factor models which characterize variation across groups at the level of a lower-dimensional subspace where the study-specific factor loadings $\Gamma_{s}$ are obtained as $\Gamma_{s}=\Lambda\Delta_{s}$ where $\Delta_{s}$ is a $d\times d_{s}$ matrix, with $d_{s}<d$, i.e. modeling the study-specific contributions as lower-dimensional perturbations of the shared factor loadings.
@@
In all of these approaches there is however a modeling limitation, i.e. they allow precisely $S$ study-specific loading matrices $\Gamma_{s}$. In practice, this assumption is often violated. We provide some examples, though a continuous spectrum of possibilities exists: (i) two or more studies exhibit high homogeneity and may share identical or nearly identical latent representations, and (ii) one study might contain highly heterogeneous subject groups, potentially forming distinct sub-populations with different latent structures (as formalized in Equation (1)). This could arise from minor experimental variations or unmeasured confounders. Additional possibilities include (iii) the existence of a subset of units from different groups that share common characteristics, or (iv) instances where some units might simultaneously exhibit characteristics of multiple groups due to various reasons.
@@
To address case (i), Grabski et al. (2023) proposed a model allowing for partially-shared latent factors using a study-specific diagonal selection matrix. The pattern of zeros and ones in this matrix determines which latent factors pertain to each specific study. The diagonal elements are then incorporated into a selection matrix governed by an Indian Buffet Process (IBP) prior (Griffiths & Ghahramani, 2011). The same work discusses a variant, where study-specific labels are unknown and inferred via mixture models. While this approach moves towards addressing case (ii), it relies on the a priori knowledge of the total number of groups into which the units are divided and to our knowledge it has not been applied yet. Case (iii) reflects the fundamental limitation of single-partition approaches in characterizing complex factor patterns. Hence, despite these recent advancements, a unifying framework that seamlessly spans from classical factor models to MSFA, encompassing all the above mentioned cases or nuanced variants, remains absent.
@@
In this paper, we propose an alternative formulation for Bayesian MSFA leveraging the concept of informed or structured sparsity (Schiavon et al., 2022, 2024; Griffin & Hoff, 2024). This approach allows for flexible modeling of varying degrees of group heterogeneity in terms of latent factors contribution. Consistently with this, we name the proposed approach Adaptive Partition Factor Analysis (APAFA). APAFA employs shrinkage priors for latent factors, using the information contained in study labels, but without enforcing a deterministic structure. This structured shrinkage is combined with increasing shrinkage priors on the number of latent factors, following a cumulative shrinkage process approach (Legramanti et al., 2020; Frühwirth-Schnatter, 2023). This construction ensures identifiable separation of the shared and specific contributions while circumventing, under suitable regularity conditions, the well-known issue of rotational ambiguity in factor loadings for the specific factors, thereby allowing straightforward posterior interpretation of these quantities. Notably, we show that the proposed method can be seen as a particular neural network model allowing us to exploit further generalizations, including the integration of subject- or study-specific continuous covariates in a flexible manner.
@@
The next section details APAFA in terms of model specification, prior distributions, and provides practical guidelines for prior elicitation. We also explore the connections to neural networks and present generalizations. In addition, we address model identifiability. Section 3 reports a comprehensive empirical assessment through simulation studies covering various data-generating scenarios. Section 4 reports two illustrative data analyses further highlighting the advantages of our approach in the motivating examples of bird species occurrence and genomics. The final discussion section explores generalizations and potential extensions. The code to reproduce our analyses available at https://github.com/elenabortolato/APAFA. The Supplementary Materials provide additional technical details on posterior computation methods, extended empirical results with further experimental analyses, and implementation specifications.
@@ 2.1 Model and prior specification
where $\Gamma=(\Gamma_{1},\dots,\Gamma_{S})$ concatenates along columns all the study-specific factor loading matrices into a $p\times k$ matrix with $k=\sum_{s=1}^{S}k_{s}$ and ${\varphi}_{i}$ is a $k$-dimensional augmented vector containing the ${\varphi}_{is}$ in Equation (1) framed with suitable pattern of zeros. Specifically, all the latent factors pertaining to group $s$ will have $\sum_{l=1}^{s-1}k_{l}$ zeros, followed by subject-specific ${\varphi}_{is}$, followed by $\sum_{l=s+1}^{S}k_{l}$ zeros. This formulation introduces two distinct factor loading matrices: $\Lambda$ containing loadings common to all units, and $\Gamma$ containing all the study-specific loadings that are activated following the non-zero patterns of $\varphi_{i}$. An illustration is depicted in Figure 1.
@@
Figure 1: Multi-study Factor model representation: the $n\times p$ data matrix $Y$ (on the left) is written as the product of the latent factor matrix H of dimension $n\times d$ by the factor loading matrix $\Lambda$ (the shared parts) plus the product of the latent factor matrix $\Phi$ of dimension $n\times k$ by the factor loading matrix $\Gamma$ (collecting all the study-specific parts), and a random noise $\epsilon$. Different shades of purple, red, and blue identify the $S=3$ studies in $Y$.
As customary in factor analysis, one can marginalize out the latent factors $\eta_{i}\sim N(0,I_{d})$, obtaining
@@
Under this characterization, one can think to have, for each $i$, a vector of dummy variables $x_{i}$ with $S$ entries characterizing the study to which unit $i$ belongs to and then assume that the $h$-th element of the vector $\varphi_{i}$ is $\varphi_{ih}=\tilde{\varphi}_{ih}\psi_{ih}$ where $\tilde{\varphi}_{ih}$ is a continuous random variable and $\psi_{ih}=f_{h}(x_{i})$ with $f_{h}$ a deterministic activation function. In De Vito et al. (2019, 2021), for example, $f_{h}(x_{i})=x_{i}^{\top}1_{S}$ where $1_{S}$ is a $S$-dimensional vector of ones. In this paper, we propose to incorporate the information contained in the $x_{i}$’s in a more flexible manner letting
@@
It is now evident that Equations (3) and (4) transform model (2) into a specific single layer neural network where the dummy variables in the $x_{i}$s are the input variables, $y_{i}$ are the $p$-dimensional output variables, $\varphi_{i}$s are the nodes of the hidden layer, $f_{h}$ are the activation functions, $\beta_{h}$ are the weights between the input and the hidden layer, and the elements in $\Gamma$ are the weights between the hidden layer and the continuous outcome $y_{i}$. The neural network representation of model (2) is reported in Figure 2. In recent years, there has been a growing interest in Bayesian analysis of neural networks. Typically, the weight parameters are assigned a prior distribution and are learned updating the prior via Bayes rule. For comprehensive reviews, see Goan & Fookes (2020), Fortuin (2022), and Arbel et al. (2023). Bayesian approaches to neural network models, however, not only apply the Bayesian inferential paradigm to learn the network parameters but also allow the neurons themselves to be stochastic (Neal, 1990; Tang & Salakhutdinov, 2013).
@@
We let the global scale parameter $\tau^{\varphi}_{h}\sim\text{Ber}(1-\rho_{h})$, where $\{\rho_{h}\}\sim\mbox{CUSP}(\alpha^{\varphi})$ follows a cumulative shrinkage process (Legramanti et al., 2020), i.e. for $h=1,2,\dots,$ $\rho_{h}=\sum_{l\leq h}w^{\varphi}_{l}$,   $w^{\varphi}_{l}=v^{\varphi}_{l}\prod_{m<l}(1-v^{\varphi}_{l})$, and $v^{\varphi}_{l}\sim\text{Beta}(1,\alpha^{\varphi})$. This approach facilitates learning the number of study-specific factors or, following the neural network interpretation of our model, the number of neurons in the hidden layer. Notably, this construction, inspired by successful applications in factor models (Legramanti et al., 2020; Schiavon et al., 2022), offers valuable insights into performing Bayesian inference for the number of nodes in the hidden layer of a Bayesian neural network, potentially connecting with existing literature on this topic (e.g Wen et al., 2016; Nalisnick et al., 2019; Fortuin, 2022; Cui et al., 2022; Jantre et al., 2023; Sell & Singh, 2023).
@@
$\tilde{\varphi}_{ih}(x_{i})\sim N(0,\tau_{h}^{\varphi}).$
Marginalizing out $\varphi_{ih}$ or $\psi_{ih}$ yelds important insights. For instance, when marginalizing out $\psi_{ih}(x_{i})$ while keeping $\tau_{h}^{\varphi}$ and $\beta_{h}$ fixed, we obtain:
@@
where $\delta_{a}$ is a Dirac mass at value $a$. Conversely, marginalizing out $\tilde{\varphi}_{ih}$ under unit variance leads to
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y_{i}\sim N(0,\Omega_{i}),\quad\Omega_{i}=\Lambda\Lambda^{\top}+\Gamma\text{diag}{\{\psi_{i}\}}\Gamma^{\top}+\Sigma.}$ (7)
Notably (6) corresponds to a simple form of conditional variational autoencoder that, differently from the standard, allows for the latent covariance structure to be informed by the independent variables. Variational autoencoders are a specific extension of neural networks, widely recognized as a descendant of classical factor models (Kingma & Welling, 2014; Sen et al., 2024).
The benefits of the proposed solution should now be evident. The shrinkage prior on the elements $\varphi_{ih}(x_{i})$ enables and promotes, yet does not mandate, the sparse representation in (2). Furthermore, it accommodates a wide range of scenarios discussed in the Introduction, including: (i) two or more studies exhibit high homogeneity and share nearly identical latent representations, (ii) some studies involve highly heterogeneous groups of subjects, potentially resulting in two or more sub-populations with distinct latent structures, or any combinations of the above. The random partition of the observation induced by the sparse structure of the $\psi_{i}$ justifies the Adaptive Partition Factor Analysis (APAFA) name adopted for the proposed solution.
@@
One remark on the activation function pertains to the potential lack of properness of the conditional posterior distribution for the parameters $\beta_{h}$, $h=1,\dots,k$. This concern is tied to the issue of perfect separation in logistic regression. Assuming the group structure encoded by the $x_{i}$ dummy variables is both accurate and relevant to the problem at hand—in other words, that the original MSFA model by De Vito et al. (2021) is correctly specified— a perfect separation in the $\psi_{ih}$ occurs with respect to the $x_{i}$ within the logistic regression. For this reason, the prior distribution on each parameter $\beta_{h}$ should be sufficiently concentrated around zero to prevent the associated conditional posterior from exhibiting a tendency toward monotonicity. To this end, we specify $\beta_{h}\sim N(0,B)$, where $B$ is a $S\times S$ diagonal matrix with elements proportional to $n^{-1}$. Thus, when perfect separation in the $\psi_{ih}$’s with respect to the $x_{is}$’s occurs, the full conditional posterior is
@@
Given the model, posterior computation is performed using Gibbs sampling, with full methodological details provided in the Supplementary Materials.